### v2

# A new Angle on Benchmarking Noncovalent Interactions

Abstract
For noncovalent interactions (NCIs), the CCSD(T) coupled cluster method is widely regarded as the ‘gold standard’. With localized orbital approximations, benchmarks for ever larger NCI complexes are being published; yet tantalizing evidence from quantum Monte Carlo (QMC) results appears to indicate that as the system size grows, CCSD(T) overbinds NCIs by progressively larger amounts, particularly when -stacking is involved. Alas, post-CCSD(T) methods like CCSDT(Q) are cost-prohibitive, which requires us to consider alternative means of estimating post-CCSD(T) contributions.
In this work, we take a step back by considering the evolution of the correlation energy with respect to the number of subunits for such -stacked sequences as acene dimers and alkadiene dimers. We show it to be almost perfectly linear, and propose the slope of the line as a probe for the behavior of a given electron correlation method. By comparison with rank-reduced CCSDT(Q) results for benzene and naphthalene dimers, we show that while CCSD(T) does slightly overbind, it does not at the level suggested by the QMC results.
keywords:
Noncovalent interactions, coupled cluster theory
1 Introduction
Theoretical predictions have become an invaluable tool in modern science. As accurate benchmarks in quantum chemistry become more and more common and numerous,1 some questions have arisen about the applicability of the underlying benchmark methods. For increasingly large systems, there are growing discrepancies2 between the interaction energies of large molecules for the main benchmark method of quantum chemistry, which is CCSD(T) (coupled cluster3 including all single and double substitutions and quasiperturbative triple excitations4) and the main benchmark method of solid state materials science, which is QMC (quantum Monte Carlo).5, 6, 7, 8.
We are trying to glance at a piece of this puzzle by extrapolating dimer interactions to larger and larger sizes using different methods. We will show that by investigating the slope of the intermolecular correlation energies with respect to monomer size, we can estimate the error of benchmark methods for much larger systems even than the ones calculated, gauging their accuracy.
Here, we will focus on CCSD(T) as the gold standard of quantum chemistry and examine all conceivable causes of a possible failure:
1.
The polarizability will increase with the system size of aromatic molecules, and so will the strength of their van der Waals interactions with each other. This polarizability may thus be increasingly hard to calculate with progressively larger basis sets. For example, even small molecules and atoms with huge polarizabilities need many diffuse functions in order to accurately describe these quantities.9
2.
In order to calculate larger systems with CCSD(T), local approximations are utilized. As the number of very small contributions to the interaction energy grows exponentially, their neglect might cumulatively result in substantial errors for larger interacting systems.
3.
While for energies, accurate CCSD(T) values are rather easily computable, CCSD(T) geometries are much scarcer, and there are very few studies actually optimizing intermolecular interaction geometries. This effect is mostly neglected, as fixed geometries are taken into account. For training or benchmarking lower-level methods like density functional theory (DFT), this a rational, or at least expected choice. Neglect of the ’relaxation energy’, however, may conceivably affect comparisons with more accurate methods.
4.
Larger intermolecular interactions are possibly overestimated by the use of perturbative triples in the (T) part, as second-order Møller-Plesset perturbation theory (MP2) strongly overestimates10 intermolecular interactions of aromatic molecules. Various papers have proffered this argument, showing that post-CCSD(T) contributions should become relevant at some point.11, 12, 13 However, one has to bear in mind that while MP2 overestimates interaction energies of stacks, MP3 overcorrects and underestimates (which has inspired the MP2.5 method14). CCSD can be seen as a summation of MP singles, doubles, and disconnected products thereof to infinite order, and underestimates interaction energies as it neglects connected triples (T), which make their first appearance in fourth order. In contrast, however, other post-Hartree-Fock or post-DFT methods such as RPA (random phase approximation15) and DFT-SAPT (density functional theory-based symmetry-adapted perturbation theory16) are argued to yield better, if not correct intermolecular interaction energies of larger dimers.11, 12
Fortunately, all these error sources can be tracked down and scrutinized for smaller systems:
1.
For intermolecular interactions with medium to large basis sets, the exact basis set limit is usually bracketed by calculations excluding and including counterpoise corrections,17, 18 with very few notable exceptions, like, for example, the formic acid dimer.19 Thus, an increasingly large gap between these two calculations (including and excluding a basis set correction) would indicate a volatility of a certain post-Hartree-Fock method. Basis set incompleteness errors can also often be estimated using localized orbital or other reduced-scaling approaches.
2.
Unlike local methods, canonical post-Hartree-Fock methods do not neglect any orbitals when calculating interaction energies.
3.
While the use of CCSD(T) gradients is very often computationally prohibitive, several structures can be considered in order to get a more complete picture of the interaction. This is usually done by just varying the distances, e.g. extending the S66 set20 of intermolecular interactions to the S66x8 set,21 which considers each complex at eight different multiples of the intermonomer separation (at fixed intramonomer geometries), for a total of 528 data points.
4.
Post-Hartree-Fock methods such as coupled cluster theory3 have a clear hierarchy. For example, when static correlation effects are present and not captured well by CCSD(T), one may attempt to walk up the cluster expansion staircase to CCSDT and CCSDT(Q), which will yield some insight about whether ‘the gold standard’ is really that golden.22, 23, 24
By addressing all of these points above, we are capable of evaluating the accuracy of different methods, including the gold standard CCSD(T), for large systems.
The new angle we propose is to compare the behavior of different correlation methods along sequences of successively expanded monomers. We shall show that not only is this behavior rather linear, but that the slopes of the lines, and specifically their deviation from the most accurate method offer a new insight into the (mis)behavior of more approximate correlation methods.
2 Computational Details
The calculations have been performed with various programs, depending on the method employed. The density functional theory (DFT) 25, 26 geometry optimizations utilizing the -B97M-V 27 functional as well as the DLPNO (Domain-Based Local Pair Natural Orbital)-CCSD(T) 28, 29 calculations have been performed with ORCA 30 , DFT-SAPT31, 32 with MOLPRO33, MP334 with QChem35 because of their RI implementation36, and all dRPA15, 37, LNO (Local Natural Orbitals)-CCSD(T)38, LNO-CCSDT(Q)39, and canonical CCSD(T)4 / CCSDT(Q)40 calculations with the MRCC41 program suite. All LNO-CCSD(T) calculations have been using their respective MP2 basis set corrections as implemented in the MRCC program. While some canonical post-CCSD(T) calculations were also carried out using MRCC, the most demanding ones were performed using a development version of the CFOUR program system42.
For the L7 (i.e., seven large noncovalent complexes) set of systems,43 geometries from the original reference have been utilized as-is. All other structures were optimized with the -B97M-V functional27 using the QZVPP44 basis set with full symmetry. For example, in the case of the benzene dimer, the D6h symmetric saddle point has been utilized. This implies that all structures are directly on top of each other, as only their interaction energy matters. They are not parallely displaced, such as the minimum structures of the benzene45 and coronene dimers46 — they are not minima at all. Despite this, the monomer geometries have either been fixed or relaxed, staying within their respective symmetry for the ethylene stacks, whereas for the acene stacks, we have employed completely relaxed geometries. The rationale for this approach is to be able to use as much symmetry as possible, since only the slope of interaction energies of increasingly large dimers matters.
Throughout this study, Dunning’s correlation consistent basis sets with and without diffuse functions, denoted cc-pVZ47 and aug-cc-pVZ48, have been put to use and will sometimes be abbreviated by DZ, TZ, QZ, and 5Z, as well as aTZ, aQZ, and a5Z. The notations {T,Q}Z and {Q,5}Z refer to basis set extrapolations from TZ and QZ or QZ and 5Z basis sets according to the familiar extrapolation formula, respectively.49
Unless specifically indicated otherwise, interaction energies have been counterpoise (CP) corrected by the standard Boys and Bernardi procedure.50
In order to shed additional light on the performance of CCSD(T)4 vs. fully iterative CCSDT51 and CCSDT(Q),40 we carried out additional calculations using two post-CCSD(T) methods that still only have CPU time scaling with system size N: CCSD(T)Λ 52, 53 and CCSDT-254, 51, 55. Cremer and coworkers56 presented a detailed perturbation theoretical analysis of the differences between full CCSDT and various approximations to it. The CCSDT-3 method omits the coupling term from the full CCSDT amplitudes equations, while CCSDT-2 additionally omits all coupling terms from the said equations. Both reduce the CPU time scaling from to at the expense of an energy difference starting at fifth order in perturbation theory:
[…]
where is the fifth-order triples-(disconnected) quadruples interaction term. CCSDT(Q) is exact to fifth order: the additional terms it introduces beyond CCSDT are
where it should be noted56 that is the Hermitian conjugate of .
[…]
3 Results and Discussion
3.1 L7 Set and the discrepancies between two reference methods
The original starting point of this study was, in fact, the discrepancy between the two reference methods CCSD(T) and FN-DMC (fixed node diffusion Monte Carlo). For this purpose, it is useful to reiterate the differences between them. In addition, we have also performed DFT-SAPT calculations for all these systems. According to the theoretical background of references 11, 12 and point one above, DFT-SAPT may be even more accurate than CCSD(T) for the interaction energies of ever larger systems. In Table 1, we show the DFT-SAPT together with the literature values of various post-Hartree-Fock methods and FN-DMC for four molecules of the L7 set43 as well as the buckycatcher.
| FN-DMC | MP3 | CCSD | DFT-SAPT | CCSD(T) | MP2.5 | MP2 | |
| Benzene2 | 10.02 | 7.220 | 6.160 | 11.861 | 11.22 | 13.420 | 19.720, 19.760 |
| GCGC | 52.02 | 36.043 | 35.760 | 54.7 | 57.02 | 56.143 | 74.743, 73.6,62 74.960 |
| C3A | 62.42 | 34.143 | 47.860 | 62.7 | 69.12 | 74.743 | 113.043, 109.2,62 109.760 |
| C2C2PD | 75.62 | 27.743 | 58.260 | 81.5 | 86.12 | 95.443 | 159.943, 154.3,62 155.760 |
| C3GC | 101.32 | 61.843 | 80.860 | 104.6 | 120.02 | 127.243 | 188.843, 182.2,62 183.760 |
| C60CPPA | 130.52 | 93.860 | 157.8 | 174.62 | 354.960 |
GCGC: guanine-cytosine tetramer; C3A: adenine…circumcoronene; C2C2PD: coronene dimer; C3GC: circumcoronene…guanine-cytosine; C60CPPA: buckminsterfullerene…buckycatcher.
It is apparent from these numbers that MP2 often overestimates the interaction energies of large conjugated molecules, whereas MP3 underestimates them. MP2.5,63 which relies on compensation between the errors of MP2 and MP3, still somewhat overestimates the interaction energies when going to larger molecules.64 The values obtained by DFT-SAPT are pretty much exactly between CCSD(T) and fixed-node diffusion Monte-Carlo (FN-DNC) values.2 If we plot interaction energies vs. number of atoms, we obtain values ranging from 0.79 (MP3) to 0.94 (FN-DMC). The slope of FN-DMC is with 1.09 kJ/mol per atom much smaller than the slope of CCSD(T) with 1.46 kJ/mol per atom- with DFT-SAPT and also MP2.5 (1.28 and 1.34 kJ/mol per atom, respectively) in this case being somewhat smaller than CCSD(T).
3.2 Intermolecular Interaction Energies of Larger Molecules
The first question is: Can we estimate these numbers also from interaction energy slopes of other molecules? In the case that they do, they may offer a hint about the accuracy of different post-Hartree-Fock methods for the interaction energies of large molecules. The systems that may come in mind are the benzene, coronene, and circumcoronene dimers, as the coronene dimer with 72 atoms was already present in the original L7 set. The parallel displaced coronene dimer is yielding a significant difference between FN-DMC with an interaction energy of 75.6 kJ/mol and CCSD(T) with an interaction energy of 86.1 kJ/mol, which is more than 10 kJ/mol. Given, however, that the circumcoronene dimer already has 144 atoms, we need other systems as touchstones for our hypothesis: that we can an interpolate the behavior for really large systems from the slopes of interaction energies of polyaromatic dimers.
To stay as close to the aromatic behavior that we see in the above sequence, we chose the acene dimer series, thus again beginning with the interaction energies of the benzene through hexacene dimers. Starting from heptacene dimer, however, the monomers will undergo cyclodimerization,65, 66, 67 which means that if we want to relax geometries like in the L7 paper (even with symmetry constraints) we will have to stop at hexacene dimer. The latter, with 82 atoms, is about the same size as coronene dimer.
As a second series, we can mix these species, and arrive at the series for the benzene-anthracene, anthracene-pentacene, and so on dimers, whereas in order to retain symmetry, there are the possibility of the CNCN, CNCN-2, CNCN-4 species, and so on – with N denoting the number of rings.
Finally, to perspective, we will also consider at ethylene stacks, as these stacks are the possible smallest conjugated carbon species which can be investigated, such as the dimers of ethylene, trans-butadiene, hexatriene, octatetraene, and so on. All these species are interesting in the context of quantum chemistry due to their extended -conjugation. As the chain length increases, the extent of -conjugation also increases, resulting in enhanced stability and stronger interactions between the molecules. For larger systems, pure London forces (as distinct from -stacking) become significant for these dimers as well.
3.3 Basis Set Effects
First, we try to evaluate the basis set effects of the different methods. If, for example, polarization would be a major issue, the dissociation energies along a chain dimer sequence would not be on a straight line, but rather be flat, as the increasing polarization would be less well described by a small number of basis functions. The number of basis functions, however, also increases linearly with system size, yielding almost perfectly straight lines for different acene series, as shown in Figure 1a-c for LNO-CCSD(T) using tight settings. As counterpoise corrections for LNO and similar methods in small molecules are problematic owing to screening out a substantial part of the ”ghost atoms” contribution [see, e.g., Ref.68], we had to exclude the first couple of cp-corrected interaction energies (for example, the benzene and anthracene dimer for the DZ basis set and the benzene, anthracene and naphthalene dimer for the TZ basis set) in each graph.
Next, we aim to evaluate the corresponding graphs for the ethylene stack series, expecting to obtain results analogous to those observed for the acene stacks.
As it becomes obvious, all of the investigated series are on almost perfectly straight lines with an value of 0.992 for all basis sets. With increasing basis set size, the lines converge nicely to slopes of 24.930.23 kJ/subunit for the acene stacks. Here, the cp-corrected slopes of the aug-cc-pVQZ, cc-pV5Z, and aug-cc-pV5Z basis sets, as well as the non-cp-corrected slopes cc-pVQZ, cc-pVQZ, and aug-cc-pV5Z basis sets fall within this range. Interestingly, while the intercepts of these graphs seem highly basis set dependent, the slopes are much less sensitive to the basis set used. For example, a basis set of aug-cc-pVTZ quality deviates in the intercept by 33% from the aug-cc-pV5Z basis set for the cp-corrected correlation energy of the acene series, but in its slope by only 3%.
For the ethylene stacks, the intercept of cc-pVTZ deviates from aug-cc-pV5Z sets by 13%, and the slope by only 3%. This trend is visible for all the systems, which implies that even quite small basis sets may give an indication about the slope (albeit not the intercept) of the correlation energy and its behavior. When comparing ethylene to acene stacks, we get about a third of the slopes of the rings (acenes) per double bond (ethylene stacks). This shows that acene stacks and ethylene stacks are still somewhat related. In contrast, the intercept of the ethylene stacks is about three times the intercept of the acene stacks. The ethylene stack series has a slope of 8.700.27 kJ if we consider all basis sets except cc-pVDZ. Based on the obtained graphs and considering basis set sizes, we conclude that utilizing the cc-pVQZ basis set is probably most advantageous for obtaining reliable slopes for both series.
3.4 Local Correlation
Having established the basis set effects, we can estimate canonical CCSD(T) at the basis set limit. Furthermore, especially with the advent of using local correlation methods and extrapolating them,69, 70, 71, 72 we can evaluate their accuracy when going to larger molecules. For this purpose, we compare these methods, especially for a small basis set of double-zeta quality. Again, all of the investigated series exhibit nearly perfect linear regression with an value of 0.996 for methods.
However, there is one caveat: When increasing the LNO cutoff settings Normal Tight VeryTight vvTight (‘very very tight’), unlike for basis sets, the slopes do not always converge to the canonical value for local CCSD(T). This makes recently promoted CPS extrapolations more erratic than, e.g., extrapolations to the complete basis set (CBS) limit. Basis sets are constructed to exhibit a monotonic convergence with increasing size to the basis set limit. In contrast, LNO and DLPNO cutoffs just define which natural orbitals are neglected- the energies do not have to increase or decrease monotonically. For example, for the cc-pVDZ basis set is applied, the differences in slopes when going from Normal to Tight to VeryTight to vvTight LNO cutoffs to canonical CCSD(T) change by 2.1, 0.8, 1.0, and 0.7 kJ/subunit for the cc-pVDZ basis set and 2.3, 0.8, 0.7, 1.3 kJ/subunit for the aug-cc-pVTZ basis set, respectively. For the latter basis set, we thus obtain a gap from the VeryTight to the vvTight setting and canonical CCSD(T). From these numbers, if we would just consider LNO-CCSD(T), we would be bound to believe that the slopes are close to converging, as we would get a slope change of 2.3, 0.8, and 0.7 kJ/subunit . This is, however, not the case, as the difference to the canonical slope is not less than 0.7 kJ/subunit .
This indicates that the CPS extrapolation, especially like in our context, should be investigated further for larger molecules, and its convergence may not be as strict as warranted.
DLPNO, while being rather close to the VeryTight and vvTight LNO-CCSD(T) cutoff, sometimes even moves into the wrong direction and is increasing its slope further when the PNO-cutoff is tightened. Interestingly, this seems to be rather basis set dependent, with the basis sets without diffuse functions (cc-pVTZ and cc-pVQZ) behaving more like the cc-pVDZ basis set. Here, an extrapolation does not work — while for the aug-cc-pVTZ basis set, the CPS extrapolation is close to the canonical CCSD(T) value.
Although the differences in slopes when going from vvTight LNO-CCSD(T) to canonical CCSD(T) are somewhat larger than e.g. when going from an aug-cc-pVTZ basis set to the CBS limit, these differences are rather small. Even for slopes of larger molecules, the errors are still minor when using LNO-CCSD(T) and/or a basis set of limited size.
For the ethylene stack series the differences in slopes when going from Normal to Tight to VeryTight to vvTight to canonical CCSD(T) change by 0.78, 0.48, -0.01, and 0.19 kJ/subunit, which also implies that the CPS extrapolation would only work for rather loose convergence criteria.
In this series, it is noteworthy that the DLPNO-CCSD(T) method with a TCutPNO value of slightly underestimates the slope but yields results that are closest to those obtained with the canonical CCSD(T) method. In contrast, the DLPNO-CCSD(T) method with TCutPNO set to overestimates the slope relative to the canonical CCSD(T), yet it still provides a more accurate slope estimate than any of the LNO methods. Furthermore, the application of CPS extrapolation produces values that are very close to those obtained with the LNO-CCSD(T) method using a vvTight threshold. However, it does not improve upon the accuracy achieved with the TCutPNO setting of .
3.5 Electronic Structure Methods
Now that we have established the slopes of canonical CCSD(T) at its basis set limit and thus its performance for large molecules, we assess the performance of other, more approximate methods.
Here, we investigate several post-Hartree-Fock methods, such as MP2, MP2.5, MP3, the dispersion contribution to DFT-SAPT, as well as dRPA. From these methods, DFT-SAPT and dRPA in particular are deemed to be as more accurate than perturbation theory for large molecules.11, 12 For the acene series, we obtain the trend of MP3 CCSD DFT-SAPT dRPA CCSD(T) MP2.5 MP2. This is hardly surprising, as MP2 very much overestimates the interaction energies of large aromatic molecules, while MP3 underestimated them, as does CCSD. Of course, the question arises about the accuracy of CCSD(T), dRPA, and DFT-SAPT.
For the ethylene stack series, we obtain the trend CCSD MP3 DFT-SAPT dRPA CCSD(T) MP2.5 MP2. As observed in the acene series, MP2 overestimates dispersion contributions while both CCSD and MP3 underestimate them. dRPA yields results similar to CCSD(T), which is clearly observed for the cc-pVZ (X=D,T,Q) basis sets. (We note73 that RPA is equivalent to a ring-diagram simplification of coupled cluster with all doubles, CCD.)
3.6 Effects of the Geometry
So far, we have used geometries for both acenes and ethylene stacks in which the distance between the two monomers were fully optimized within its (high) symmetry. In order to show that the linear behavior is not an artefact, it may be worth to investigate another geometry with a different intermonomer distance. Here, we chose the stacks at the rather large optimized distance of the ethylene dimer, as shorter distances (e.g. at the hexatetraene distance) would result in a positive interaction energy for the ethylene dimer. The ethylene stack series with relaxed geometry exhibits a slope of 8.70 0.27 kJ/subunit. In contrast, the same series with the fixed geometry at the minimum ethylene dimer distance of 4.46 Å shows a much smaller slope of 1.26 0.24 kJ/subunit, where the uncertainty is about one-fifth of the actual value.
Results obtained using more approximate methods are identical for both types of geometries, with the only exception being that the slope determined by dRPA is closer to the slope obtained by CCSD(T) in the case of fixed geometries.
All methods analyzed in this study demonstrate a better convergence of the LNO approaches compared to canonical CCSD(T) for relaxed geometries, irrespective of the basis set used. This is attributed to the fact that for these geometries, the systems energy is closer to the true potential energy minimum for the respective method and basis set. This conclusion is supported by the observation that the absolute values of total energies and correlation energies are greater. Similarly, basis set convergence is more consistent for relaxed geometries, as the electron distribution and correlation effects are more accurately aligned with the optimized structure, resulting in more consistent convergence behavior as the basis set is expanded.
3.7 Best Estimates of Slopes
| Method | |||||||||
| CCSD(T) | post-CCSD(T) | ||||||||
| Slopes | Basis Set | MP2 | CCSD | Tight | VeryTight | canonical | CCSDT-2 | CCSDT | CCSDT(Q) |
| 12 | DZ | -24.19 | -13.84 | N/A | N/A | -16.67 | -16.17 | -15.92 | -16.32 |
| 15 | DZ | -25.56 | -13.94 | -19.43 | -19.84 | -16.93 | -16.43 | -16.18 | -16.58 |
| 12 | {T,Q}Z | -32.43 | -18.46 | N/A | N/A | -23.14 | -22.44∗ | -22.09∗ | -22.65∗ |
| 15 | {T,Q}Z | -33.91 | -18.72 | -24.89 | -23.40 | -22.71∗ | -22.35∗ | -22.97∗ | |
| 16 | {T,Q}Z | -33.97 | -18.78 | -24.95 | -23.46 | -22.77∗ | -22.41∗ | -22.98∗ | |
| 13 | a{Q,5}Z | -23.79 | -23.36 | ||||||
| 16 | a{Q,5}Z | -33.58 | -18.39 | -24.55 | -24.12 | -23.07 | -22.37∗ | -22.02∗ | -22.58∗ |
∗ The post-CCSD(T) slopes have been scaled by a factor of 1.4 to account for the small cc-pVDZ basis set size — see discussion below and Supporting Information.
| Method | ||||||||
| CCSD(T) | post-CCSD(T) | |||||||
| Slopes | Basis Set | MP2 | CCSD | Tight | VeryTight | canonical | CCSDT | CCSDT(Q) |
| 12 | DZ | -27.16 | -15.93 | -20.95 | -20.12 | -19.32 | -18.54 | -18.94 |
| 15 | DZ | -28.53 | -16.04 | -21.71 | -20.41 | -19.49 | -18.71 | -19.11 |
| 12 | {T,Q}Z | -32.66 | -18.40 | -24.38 | -23.60 | -23.09 | -22.00∗ | -22.56∗ |
| 15 | {T,Q}Z | -33.92 | -18.57 | -24.89 | -23.26 | -22.17∗ | -22.73∗ | |
| 16 | {T,Q}Z | -33.95 | -18.60 | -24.92 | -23.29 | -22.20∗ | -22.76∗ | |
| 12 | a{Q,5}Z | -24.09 | -23.00 | -22.48 | -21.39∗ | -21.95∗ | ||
| 13 | a{Q,5}Z | -24.28 | -23.22 | -22.71 | -21.61∗ | -22.17∗ | ||
| 16 | a{Q,5}Z | -33.47 | -18.46 | -24.44 | -23.39 | -22.87 | -21.78∗ | -22.34∗ |
∗ The post-CCSD(T) slopes have been scaled by a factor of 1.4 to account for the small cc-pVDZ basis set size — see discussion below and Supporting Information.
Usually, the best estimates in energies are obtained by going to ever larger basis sets and then just extrapolating the energies to the complete basis set (CBS) limit. In our case, however, we can either calculate ever larger systems along the sequences studied with a smaller basis set or increase the basis set size thus adding one dimension to the problem. Our results are summarized in Tables 2 and 3. The slopes obtained from just the first two acene dimers, namely benzene2 and naphthalene2, barely differ compared to those of the whole series up to pentacene; the cp-corrected DZ CCSD(T) slope changes by 0.26 kJ/subunit, and the cp-uncorrected DZ CCSD(T) slope even less at 0.17 kJ/subunit. In general, including larger systems in the fits increases the slopes. From these values, including the cc-pV{T,Q}Z basis set extrapolation which is still feasible for canonical CCSD(T) for the benzene and naphthalene dimer, we can estimate the canonical CCSD(T)/cbs values for the benzene-pentacene series, which will be -23.14-0.26=-23.40 kJ/subunit for the cp-corrected CCSD(T) results and -23.09-0.17=-23.26 kJ/subunit for cp-uncorrected CCSD(T). Finally, this values can then be extended to the hexacene dimer with an aug-cc-pV{Q,5}Z basis set extrapolation.
Of more interest are the post-CCSD(T) slopes. The CCSDT and CCSDT(Q) results for acene dimers (benzene and naphthalene dimers) included in Tables 2 and 3 were obtained using rank-reduced coupled cluster methods described in Refs. 74, 75, 76. In these calculations, the coupled cluster equations are solved within a certain excitation subspace instead of the full space as in the canonical methods. This allows for reducing the costs of the calculations, but introduces an additional parameter (size of the excitation subspace) as a parameter in the rank-reduced formalism. Therefore, it is important that the results are stable with respect to the adopted value of this parameter. In the Supporting Information, we provide a detailed analysis of the convergence of the rank-reduced results with respect to size of the excitation subspace for benzene dimer and naphthalene dimer, showing that the numerical values reported here are sufficiently well converged for the present purposes. The only additional approximation employed in the rank-reduced calculations is the Cholesky decomposition of two-electron integrals. We used the full pivoting variant of this decomposition with a VeryTight threshold () for the diagonal elements. The errors resulting from the decomposition with these settings impact the calculated interaction energies by no more than a few thousands of kJ/mol. The CCSDT method reduces the slopes (in absolute value) by a sizeable 0.75 (cp-corrected DZ) and 0.78 (standard DZ) kJ/subunit. However, CCSDT(Q) (again, in absolute value) reverts this by 0.4 in both instances, shrinking the CCSD(T) to CCSDT(Q) differences in the slopes to 0.35 (cp-corrected DZ) and 0.38 (uncorrected DZ) kJ/subunit.
(We note in passing that the (Q) contribution found here for benzene dimer, 0.50 kJ/mol, is just over half of the value of Karton and Martin77, 0.187–0.200 kcal/mol, or 0.78–0.84 kJ/mol. The latter was obtained from a thermochemical cycle combined with a CCSDT(Q) calculation in an unpolarized double-zeta basis set, and serves as a cautionary tale in this regard.)
For the Hartree-Fock exchange part, this is rather simple, as all calculations were performed at the largest basis sets and the cp-corrected and cp-uncorrected slopes are 11.95 and 11.96 kJ/subunit, respectively, yielding exchange-correlation slopes of -10.5 0.2 kJ/mol per acene ring dimer. For the largest acene dimer hexacene, this would make an interaction energy of approximately 63.0 kJ/mol, considering our best estimates. This value, however, is neglecting the intercepts of 3.3 for Hartree-Fock exchange and 1.9 for correlation, arriving at 57.8 1.4 kJ/mol. CCSD(T), for comparison, would yield slightly larger slopes of -11.1 and -10.9, with an interaction energy of 66.0 for the hexacene dimer. Including the intercepts (3.3 and 2.0), we arrive at a final CCSD(T) energy of 60.7 1.5 kJ/mol.
| Method | |||||||||
| CCSD(T) | post-CCSD(T) | ||||||||
| Slopes | Basis Set | MP2 | CCSD | VeryTight | vvTight | canonical | CCSDT-2 | CCSDT | CCSDT(Q) |
| 12 | DZ | -6.53 | -4.74 | -5.56 | -5.58 | -5.47 | -5.35 | -5.36 | -5.49 |
| 13 | DZ | -6.99 | -4.89 | -5.83 | -5.81 | -5.69 | -5.56 | -5.54 | -5.67 |
| 14 | DZ | -7.44 | -5.08 | -6.12 | -6.07 | -5.95 | -5.81 | -5.79 | -5.92 |
| 12 | TZ | -8.40 | -5.82 | -7.01 | -7.05 | -6.95 | -6.81∗ | -6.80∗ | -6.93∗ |
| 17 | TZ | -9.93 | -6.38 | -7.95 | -7.96 | -7.77 | -7.62∗ | -7.61∗ | -7.74∗ |
| 12 | QZ | -9.06 | -6.24 | -7.63 | -7.67 | -7.54 | -7.40∗ | -7.39∗ | -7.52∗ |
| 15 | QZ | -10.33 | -6.72 | -8.40 | -8.42 | -8.24 | -8.10∗ | -8.08∗ | -8.21∗ |
| 17 | QZ | -8.54 | -8.58 | -8.37 | -8.23∗ | -8.21∗ | -8.35∗ | ||
| 12 | a{T,Q}Z | -9.53 | -6.55 | -8.04 | -8.07 | -7.98 | -7.84∗ | -7.82∗ | -7.96∗ |
| 17 | a{T,Q}Z | -8.79 | -8.65∗ | -8.64∗ | -8.77∗ | ||||
| 12 | a{Q,5}Z | -9.54 | -6.55 | -8.07 | -8.04 | -7.98 | -7.84∗ | -7.83∗ | -7.96∗ |
| 15 | a{Q,5}Z | -8.67 | -8.53∗ | -8.52∗ | -8.65∗ | ||||
| 17 | a{Q,5}Z | -8.79 | -8.65∗ | -8.64∗ | -8.77∗ | ||||
∗ The post-CCSD(T) slopes have been scaled by a factor of 1.4 to account for the small cc-pVDZ basis set size, see discussion below and Supporting Information.
| Method | |||||||||
| CCSD(T) | post-CCSD(T) | ||||||||
| Slopes | Basis Set | MP2 | CCSD | VeryTight | vvTight | canonical | CCSDT-2 | CCSDT | CCSDT(Q) |
| 12 | DZ | -7.03 | -4.87 | -5.91 | -5.93 | -5.72 | -5.59 | -5.59 | -5.72 |
| 13 | DZ | -7.60 | -5.15 | -6.26 | -6.24 | -6.09 | -5.94 | -5.92 | -6.06 |
| 14 | DZ | -8.14 | -5.44 | -6.66 | -6.62 | -6.46 | -6.30 | -6.28 | -6.41 |
| 12 | TZ | -9.09 | -6.27 | -7.61 | -7.64 | -7.53 | -7.38∗ | -7.36∗ | -7.49∗ |
| 17 | TZ | -10.71 | -6.94 | -8.62 | -8.62 | -8.46 | -8.30∗ | -8.28∗ | -8.42∗ |
| 12 | QZ | -9.41 | -6.45 | -7.91 | -7.93 | -7.81 | -7.65∗ | -7.63∗ | -7.76∗ |
| 15 | QZ | -10.72 | -6.97 | -8.71 | -8.72 | -8.55 | -8.39∗ | -8.37∗ | -8.51∗ |
| 17 | QZ | -8.83 | -8.85 | -8.67 | -8.51∗ | -8.49∗ | -8.63∗ | ||
| 12 | a{T,Q}Z | -9.44 | -6.30 | -7.78 | -7.79 | -7.72 | -7.56∗ | -7.54∗ | -7.68∗ |
| 17 | a{T,Q}Z | -8.65 | -8.49∗ | -8.47∗ | -8.60∗ | ||||
| 12 | a{Q,5}Z | -9.58 | -6.53 | -8.04 | -8.04 | -7.95 | -7.80∗ | -7.78∗ | -7.91∗ |
| 15 | a{Q,5}Z | -8.69 | -8.54∗ | -8.52∗ | -8.65∗ | ||||
| 17 | a{Q,5}Z | -8.88 | -8.72∗ | -8.70∗ | -8.84∗ | ||||
∗ The post-CCSD(T) slopes have been scaled by a factor of 1.4 to account for the small cc-pVDZ basis set size, see discussion below and Supporting Information.
An analogous estimation is performed for the ethylene stack series in Table 4 and 5. To extrapolate CCSDT and CCSDT(Q) values, we utilize the difference between CCSDT and CCSD(T) or between CCSDT(Q) and CCSD(T) results obtained with the cc-pVDZ basis set. The most accurate CCSD(T) calculations were performed using the cc-pVTZ basis set for the entire ethylene stack series (ethylene, tetraene, pentaene up to tetradecaneheptaene) and the cc-pVQZ basis set for ethylene - decanepentaene. These results were subsequently used to extrapolate values to larger basis sets. Interestingly, canonical CCSD(T) is even slightly underestimating the correlation interaction energy slopes relative to CCSDT(Q), showing that the correlation slopes obtained are rather dependent on the systems chosen. Overall, the post-CCSD(T) contributions are also smaller for the ethylene stacks, as the CCSDT-CCSD(T) difference amounts to 2.7% of the correlation contribution to the interaction energy. For the acene series, the CCSDT-CCSD(T) difference is more substantial at 4.6% — presumably because of the aromatic rings.
The more economical CCSD(T)Λ method yields qualitatively the same answer as CCSDT-3 and CCSDT-2, but not quantitatively.
3.8 Assessment of the results when extrapolating to larger molecules
How would this translate to other, similar systems and their intermolecular interactions, such as the coronene dimer from the L7 set of molecules and the buckycatcher? There are several aspects to this: Clearly, the interaction energy of the buckycatcher of approximately 175 kJ/mol at the CCSD(T) level of theory is much larger than the interaction energy of the hexacene dimer with our best estimate at approximately 61 kJ/mol. The latter, with six rings, is more comparable to the coronene dimer with seven rings, and an interaction energy of 76 kJ/mol (FN-DMC) to 86 kJ/mol (CCSD(T)). However, the small discrepancy between CCSD(T) and CCSDT(Q) is rather telling, as it is only 1.5 kJ/mol for the hexacene dimer. Even extrapolating the same slope to the buckycatcher, the deviation would only sum up to approximately 4 kJ/mol difference, indicating that the deviations between CCSDT(Q) and CCSD(T) for intermolecular interactions are an order of magnitude smaller than the 44 kJ/mol difference between FN-DMC and CCSD(T) of Table 1.
If we do not compare interaction energies, but rather MP2-MP3 and MP2-CCSD differences of 128 (MP2-MP3) and 98 kJ/mol (MP2-CCSD) for the coronene dimer, and a (rather large) MP2-CCSD difference of 261.1 kJ/mol for the buckycatcher, we will arrive at different numbers by extrapolating to these values, considering the slopes obtained above for the acenes. This MP2-CCSD difference of the slope is 15.3 0.1 kJ/mol per acene ring at the basis set limit, taking cp-corrected and uncorrected values into account. Interestingly, the MP2-CCSD difference is highly basis set dependent, as for the cc-pVDZ basis set, the difference in the slope is only 10.8 0.5 kJ/mol per acene ring. The deviations in the slopes between CCSD(T) and CCSDT(Q) are, in comparison, only 0.36 0.02 kJ/mol, which is a factor of 40 smaller than the MP2-CCSD difference.
Extrapolating this difference again to the buckycatcher, this would amount to a deviation of 5.6 kJ/mol (and 2.3 kJ/mol for the coronene dimer). Unfortunately, we have CCSDT(Q) values using only a double-zeta basis set. However, we can compute other approximate methods using a larger basis set, compare, and then scale the CCSDT(Q)-CCSD(T) differences in the slopes by these values. For the acenes, the MP2-CCSD(T) slopes differ by 21% when going from a double-zeta basis set to an extrapolated cc-pV{T,Q}Z value . In contrast, the CCSD-CCSD(T) slopes differ by 57% for the same basis sets (DZ and cc-pV{T,Q}Z). As we can see from Table 2, the CCSDT-2 slopes are rather close to CCSDT(Q), and the CCSDT-2-CCSD(T) slopes behave slightly more like CCSD-CCSD(T) for a small basis set of triple-zeta quality, see Supporting Information. Because of this, we assume that at the double-zeta basis set for the CCSD(T)-CCSDT(Q) slope the difference is underestimated by a maximum of 40%. Hence, we would obtain a deviation from CCSD(T) of 7.9 kJ/mol for the buckycatcher and 3.2 kJ/mol for the coronene dimer.
These values of CCSDT(Q) underestimating CCSD(T) by 8 and 3 kJ/mol are much smaller than the discrepancies of 44 and 10 kJ/mol for the buckycatcher and coronene dimer found in between FN-DMC and CCSD(T), respectively. Hence, our estimates are much closer to CCSD(T) than to FN-DMC (when compared to Table 1).
The large deviations in the slopes of CCSD and MP2 in comparison to that of CCSD(T) and CCSDT(Q) also imply that a different angle when looking at intermolecular interactions is needed, and approximate methods may unfortunately not provide a full picture.
For the L7 systems and the buckycatcher dimer in question, the ethylene stacks are not as much of interest, as their post-CCSD(T) effects can be considered as negligible here, the CCSD(T) and CCSDT(Q) slopes are extremely similar, arriving at the CCSDT(Q) line y = (-8.7 0.2 kJ/mol)x + 6.5 0.2 kJ/mol per ethylene double bond. In comparison, the final best CCSDT(Q) lines obtained for the acenes are = (-10.5 0.2 kJ/mol) + 5.2 0.3 kJ/mol for the full set of the six acenes. The interaction energies of benzene () and naphthalene () sandwich structures are -6.5 0.2 kJ/mol and -15.5 0.2 kJ/mol, respectively: the resulting slope would be slightly different from the one discussed above because the two data points include the Hartree-Fock association energy, which has its own slope. The interaction energies of ethylene and trans-butadiene, within the geometric constraints applied, are -1.7 kJ/mol and -7.3 kJ/mol. These small molecules or the line of the series can serve as reference for other methods when testing the increase in interaction energy for such species.
4 Conclusions
In case we move to ever larger molecules, we propose a new angle of looking at intermolecular interactions of molecular dimers. Although local methods already provide a good, alternative avenue for calculating these, we are unfortunately still away from drawing definite conclusions about the accuracy of electronic structure methods, such as the gold CCSD(T) standard.
Here, we use different dimer series exhibiting very linear increases of the correlation energy, such as acene and ethylene stacks. We were able to show that none of the following factors disrupt this linear behavior: the choice of basis set; the local correlation cutoffs; the electronic structure method; fixed vs. relaxed geometry. This is obvious from all figures in this paper.
The effects of these factors (basis set, local correlation, electronic structure method, geometry) on the slopes are however sizeable.
Concerning the basis sets, cc-pVTZ already is sufficient to obtain acceptable slopes, with a deviation in the slopes of 1.5 kJ/mol per ring for the acene series. For local correlation methods, only the tightest cutoffs (vvTight or TCutPNO=7) achieve this accuracy, whereas other cutoffs can deviate quite significantly with more than 4 kJ/mol per ring.
Finally, we are able to estimate the interaction of large acenes and ethylene stacks which are of similar interaction energy size or MP2-CCSD difference than the L7 set of molecules by comparing CCSD(T) to CCSDT(Q) slopes: Here, the deviations are much smaller than anticipated, with the CCSDT(Q)-CCSD(T) difference in the slope of 0.4 kJ/mol per ring. For comparison, this value is smaller than the deviation of canonical CCSD(T) from local CCSD(T) with the tightest cutoffs or medium-size basis sets of (aug-)TZ quality from the basis set limit, showing that these two effects can still pose major challenges.
By converging all these effects on the slopes to 0.2 kJ/mol per ring or double bond, we can estimate deviations of interaction energies from CCSD(T) at the basis set limit for large acene and ethylene stacks, and ultimately very large molecules including those which are having huge polarizabilities.
In case of the acene dimer series, CCSD(T) exhibits an increased interaction energy of 3.4% in comparison to CCSDT(Q), which is sizeable, but nowhere near the level that would explain the discrepancies to FN-DMC. For the ethylene stacks, on the other hand, there is no notable difference between CCSD(T) and CCSDT(Q). For the ethylene stacks, on the other hand, there is no notable difference between CCSD(T) and CCSDT(Q). Thus, with some limitations, we can reconfirm the ”gold standard” of basis set limit CCSD(T) for intermolecular interactions.
Acknowledgments
This research was supported by an internal grant to JMLM from the Uriel Arnon Memorial Fund for Artificial Intelligence in Materials Research, as well as by a very generous allocation of CPU time on the Weizmann Faculty of Chemistry HPC facility chemfarm. The authors thank its administrators, Dr. Mark Vilensy and Andrei Vasilev, for their kind assistance. ADB was supported by the Weizmann Institute of Science as a Weston Visiting Scholar.
PDF document with full raw data, including additional discussions, extra methods used, data for further polyaromatic hydrocarbons, can be found in the Supporting Information.
Cartesian coordinate (xyz) files of the structures used in this paper are given.